media%2F0b5%2F0b52bdd8-e722-48ff-8610-44
Answer #1


arms streched angular omentum L1 = I1*w1 = 100*2 = 200

arms tucked L2 = I28w2 = 75*w2

from momentum conservation

L2 = L1


75*w2 = 200


w2 = 2.67 rps